OpenStudy (gabylovesyou):

Assuming x > 0, which of these expressions is equivalent to 11 sqrt(245x^3) + 9 sqrt(45x^3) ?

OpenStudy (gabylovesyou):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

hmm I'm assuming there are answer choices? if not, then just try to simplify 11 sqrt(245x^3) + 9 sqrt(45x^3) as much as possible

OpenStudy (gabylovesyou):

what comes after 104 sqrt(5) sqrt(x^3)

jimthompson5910 (jim_thompson5910):

how are you getting that

OpenStudy (gabylovesyou):

idk how to do the sqrt thing so im trying my best lol

OpenStudy (gabylovesyou):

ok forget what i said how do i simplify that ?

jimthompson5910 (jim_thompson5910):

can you factor 245 into two numbers where one factor is a perfect square?

jimthompson5910 (jim_thompson5910):

yes? or no?

OpenStudy (gabylovesyou):

i can try..

jimthompson5910 (jim_thompson5910):

alright tell me what you get

OpenStudy (gabylovesyou):

5 * 49 ?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so \[\large \sqrt{245} = \sqrt{5*49}\] \[\large \sqrt{245} = \sqrt{49*5}\] \[\large \sqrt{245} = \sqrt{49}*\sqrt{5}\] \[\large \sqrt{245} = 7*\sqrt{5}\] see how I'm getting this?

OpenStudy (gabylovesyou):

yes i got it

jimthompson5910 (jim_thompson5910):

ok so this means \[\large \sqrt{245x^3} = \sqrt{5*49x^2*x}\] \[\large \sqrt{245x^3} = \sqrt{49x^2*5x}\] \[\large \sqrt{245x^3} = \sqrt{49x^2}*\sqrt{5x}\] \[\large \sqrt{245x^3} = 7x*\sqrt{5x}\]

jimthompson5910 (jim_thompson5910):

Then we can cay \[\large 11\sqrt{245x^3} = 11\sqrt{5*49x^2*x}\] \[\large 11\sqrt{245x^3} = 11\sqrt{49x^2*5x}\] \[\large 11\sqrt{245x^3} = 11\sqrt{49x^2}*\sqrt{5x}\] \[\large 11\sqrt{245x^3} = 11*7x*\sqrt{5x}\] \[\large 11\sqrt{245x^3} = 77x*\sqrt{5x}\]

jimthompson5910 (jim_thompson5910):

Use these ideas to simplify \(\large 9\sqrt{45x^3}\)

OpenStudy (gabylovesyou):

i have to find what times what = 45 right ?

jimthompson5910 (jim_thompson5910):

yep, factor it and make sure one factor is a perfect square

OpenStudy (gabylovesyou):

my OS is acting weird sorry....... can it be 9 * 5 ?

OpenStudy (gabylovesyou):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

yep, 45 = 9*5 use this to get \[\large 9\sqrt{45x^3} = 9\sqrt{9*5x^2*x}\] \[\large 9\sqrt{45x^3} = 9\sqrt{9x^2*5x}\] \[\large 9\sqrt{45x^3} = 9\sqrt{9x^2}*\sqrt{5x}\] \[\large 9\sqrt{45x^3} = 9*3x*\sqrt{5x}\] \[\large 9\sqrt{45x^3} = 27x*\sqrt{5x}\]

OpenStudy (gabylovesyou):

@jim_thompson5910 what happens after that ?

jimthompson5910 (jim_thompson5910):

\[\large 11\sqrt{245x^3} + 9\sqrt{45x^3} \] turns into \[\large 77x*\sqrt{5x} + 27x*\sqrt{5x} \]

jimthompson5910 (jim_thompson5910):

what's next

OpenStudy (gabylovesyou):

ummmmmmm

OpenStudy (gabylovesyou):

add 77 + 27 ?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

to get ?

OpenStudy (gabylovesyou):

104x

OpenStudy (gabylovesyou):

then just add the sort(5x) .... so the answer would be 104x sqrt(5x)

jimthompson5910 (jim_thompson5910):

so \[\large 77x*\sqrt{5x} + 27x*\sqrt{5x} \] turns into \[\large 104x*\sqrt{5x} \] yep you got it

OpenStudy (gabylovesyou):

thank you :)

jimthompson5910 (jim_thompson5910):

sure thing